OpenStudy (anonymous):

Find the constant of variation k for the inverse variation. Then write an equation for the inverse variation. y = 5 when x = 7.5

OpenStudy (jhannybean):

An inverse equation follows the formula \[\large y=\frac kx\]You're given \[\large y=5 \ , \ x= 7.5\] plug these values into your equation. \[\large 5 = \frac{k}{7.5}\] solve for k. \[\large k = 5(7.5)\]\[\large k = 37.5\] rewrite equation \[\large y= \frac{37.5}{x}\] and now we can check to see if our values are correct \[\large 5 = \frac{37.5}{7.5}\]\(\large 5 = 5\) \(\large \checkmark\)